OpenStudy (anonymous):

|8t-9| less than or equal to 6 Find the solution and answer in interval notation. Please explain how to do it, I'm really confused!

OpenStudy (anonymous):

\[|8t-9| \le 6\]

OpenStudy (helder_edwin):

\[ \large |a|\leq b\Leftrightarrow\begin{cases} b\geq0\\\text{and}\\ -b\leq a\leq b \end{cases} \]

OpenStudy (anonymous):

Can you explain? I haven't seen that before

OpenStudy (helder_edwin):

\[ \large |8t-9|\leq6 \] becomes \[ \large -6\leq 8t-9\leq 6 \]

OpenStudy (anonymous):

Alrighty, and then what do you do? o:

OpenStudy (helder_edwin):

first add 9 to everything then divide by 8

OpenStudy (anonymous):

Oh okay I got it! Thank you!

OpenStudy (helder_edwin):

u r welcome

OpenStudy (anas.p):

@helder_edwin the question has mod in it, dont you have to square it to remove the mod I dont know how you did it but please tell me what you thin of this method?? \[\left| 8t-9 \right|\le6\]\[(8t-9)^{2}\le(6)^{2}\]\[(64t^{2}-144t+81)\le36\]\[64t^{2}-144t+45\le0\]I am not sure what to do next, but shouldn't it start like this??

OpenStudy (helder_edwin):

@Anas.P

OpenStudy (helder_edwin):

@Anas.P r u there?

OpenStudy (helder_edwin):

well. what u did u fine. but see what Insist did: \[ \large |8t-9|\leq6 \] \[ \large -6\leq 8t-9\leq6 \] \[ \large 3\leq 8t\leq15 \] \[ \large \frac{3}{8}\leq t\leq\frac{15}{8} \] so the solution is \([3/8,15/8]\).

OpenStudy (helder_edwin):

whereas what u did is incomplete. u haven't solved the inequation just yet.

OpenStudy (helder_edwin):

the way Insist solved the inequation is easier.

OpenStudy (helder_edwin):

to finish what u started: now factor the quadratic polynomial u got: \[ \large 64t^2-144t+45\leq0 \] \[ \large (8x-15)(8x-3)\leq0 \] from this u get the points \[ \large x=3/8\qquad\vee\qquad x=15/8 \]

OpenStudy (anas.p):

hey thanks.... gotya... nice way of solving, Thanks a lot..

OpenStudy (helder_edwin):